OpenStudy (anonymous):

Can someone Help Me With 2 Algebra questions? ;( Please please , ill give medals and fan :( <3 pleaseeeeeeee! thank yous!:3

OpenStudy (anonymous):

Yah of course post the questions!

OpenStudy (anonymous):

I Need Help PLEASE :'( Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 |2x - 3| = -17 Equation #2 |5x + 3| = 12 -Equation #1 and Equation #2 have the same number of solutions. - Equation #1 has more solutions than equation #2. -Equation #1 has fewer solutions than equation #2. -None of the statements above describe the number of solutions to equations shown.

OpenStudy (anonymous):

i ve already answered it both the equations have same number of solutions

OpenStudy (anonymous):

it didnt say you did? :( im sorry but i have one more

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation#2 4 - 3x = 7x - 6 8 + 4x = -10 + x The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describe solutions to equations shown.

OpenStudy (anonymous):

The solution to equation #1 is larger than the solution to equation #2. because the solution for equation 1 is 1 solution for equation 2 is -6

OpenStudy (anonymous):

thank you so much!